A monopolist faces the following demand curve:

Q = 80 – 0.2P

Where Q is the weekly production and P is the price, measured in $/unit. The firm’s cost function is given by C = 100 + 20Q2 . Assuming the firm maximizes profits,

  1. Find the equation describing the marginal revenue (MR) curve.
  2. What is the level of production (Q), price (P), and total profit (π) per week?
  3. If the government decides to levy a per-unit tax of 50 $/unit on this product, what will be the new level of production (Q), price (P), and total profit (π)?
Answer #1

=80-0.ap Domand: 7 P = 400 - 59 2 P. last function : C 5100 +208 Total Revenue = Th (400-52):9 Marginal Revenue = aCTR) 400 -Simu gort s Putting sot a per-unit tax of Hence cost function will now become on this product c=100 + 20 Q2 + 50 Q 190 = 400